# Detecting Harmful Memes with Decoupled Understanding and Guided CoT Reasoning

### v1

@@ C.1 Visual Information Extraction
Politicians
(i) Is any politician or celebrity portrayed in the image? If yes, who?[Fm] (i) Is any head of state portrayed in the image?[Fm] (i) Is Donald Trump depicted in the image?[Fm] (i) Is Joe Biden depicted in the image?[Fm] (i) Is Barack Obama depicted in the image?[Fm] (i) Is Hillary Clinton depicted in the image?[Fm] (i) Is Bernie Sanders depicted in the image?[Fm] (i) Is Gary Johnson depicted in the image?[Fm] (i) Does this image feature Joe Biden and Barack Obama?[Fm]
Political Issues
(i) Is any political party explicitly involved in this image?[Fm] (i) Is LGBTQ+ community or LGBTQ+ individual portrayed in this image?[Fm] (i) Is any individual of Middle Eastern descent portrayed in the image?[Fm] (i) Is any protected racial or minority group (such as African Americans or other colored people) portrayed in this image?[Fm]
@@
Describe
(i) FHM: What is shown in the image? Describe within two sentences, ignoring any overlaid text or caption. (i) HarMeme and Harm-P: What is shown in the meme? (i) MultiOff: What is shown in this image? (i) MAMI: The overlaid text on the image reads:[OCR]. Question: What is shown in the image? Describe within three sentences. DO NOT assume the nature of the image’s tone or intent as humorous, comical, playful or lighthearted in your description. (i) PrideMM: This is an online meme related to LGBTQ+ pride movement. What is this meme about? Note: DO NOT ASSUME the nature of the meme’s tone and intent as humorous or lighthearted. Describe in a neutral tone.
C.2 Information Integration
After acquiring visual information leveraging LMMs, we prompt unimodal text-only LLMs to integrate the extracted visual cues (along with the overlaid caption [OCR] for some datasets) into a coherent, unified description of the content of each meme. We experiment with four open-source small-scale LLMs with no more than 14 billion parameters i.e., Qwen2.5-14B Qwen et al. (2025), Mistral-12B Mistral (2024), Qwen2.5-7B Qwen et al. (2025) and Llama3.1-8B Grattafiori et al. (2024). All LLMs are the instruction-tuned versions. The visual information gathered is denoted by [VIG]. Examples of integration prompt templates:
FHM
Given the following information provided about an image, and disregarding any information about overlaid text or captions, synthesize and rephrase these information into a coherent and unified description of the image’s content. Information:[VIG]
PrideMM
Given the following information provided about an online meme, synthesize and rephrase these information into a unified, coherent, and neutral description of the meme content. DO NOT mention the meme’s tone and intent as humorous or light-hearted in the description. Information:[VIG]. The overlaid caption or text recognized in the meme reads:[OCR]
Section C.2 converts memes to text. The following sections (including the proposed guided CoT reasoning) are based on the generated meme description, denoted by [M2T]. Chain-of-Thought prompting instruction (e.g., “Now, let’s think step by step:”, “Now, let’s analyze by applying all the guidelines one by one:”) is denoted by [CoT].
C.3 Fine-grained Target Identification
@@
Protected Group Detection
Given the following image-caption content which may or may not be an online meme, analyze: what vulnerable protected group(s) might be involved in the content? Here are some guidelines for your reference: A. If the content does not appear to involve any specific protected group, just output“No specific protected group involved.” B. If the image content does involve specific protected group(s), choose your answer(s) from the specified list (you may choose multiple options if there are more than one protected groups involved): 1. Women (Female); 2. LGBTQ Community; 3. People with Disabilities; 4. Muslims and Islamic culture; 5. Individuals of Middle Eastern descent; 6. Jewish individuals; 7. Individuals of African descent; 8. African Americans; 9. Individuals of East Asian descent; 10. Individuals of South Asian descent; 11. Native Americans; 12. Other protected groups not listed. Here is the image-caption content you need to analyze:[M2T]. The caption overlaid on the image reads[OCR].[CoT]
We denote the target protected groups identified as $TG_{1}$, $TG_{2}$,…,$TG_{n}$. The following template prompts LLMs with few-shot human-defined hateful examples [FS] to sequentially generate more hateful forms targeting each protected group:
Hateful Forms Generation
1. Provide examples of commonly found harmful stereotypes and forms of offensive, hateful content against $TG_{1}$ in online memes. Provide only phrases or terms without detailed explanations e.g.,[FS]. 2. Provide examples of commonly found …against $TG_{2}$
@@
Entity Identification
Given the following description of an online meme related to LGBTQ+ movements, { (i) Country and Region: does the meme explicitly reference any country or region where LGBTQ+ identities or advocacy are discouraged? (i) Politics: does the meme explicitly involve or mention politicians, political figures, political parties, ideologies, or groups? (i) Company: does the meme explicitly touch on topics about corporate involvement in LGBTQ+ movements? (i) Specific Individual: analyze: Does the meme content involve any specific individual? <Definition of Specific Individual> (i) Organization: analyze: Does the meme content address or discuss organizational involvement related to LGBTQ+ issues? <Definition of Organization> (i) LGBTQ+ Subgroup: analyze: what specific subgroup(s) within the LGBTQ+ community is/are particularly mentioned, referenced or implicitly implied in the meme content? <Subgroup Identification Guidelines> }**Description of the meme content**: [M2T]{[Fm]|[CoT]}
Target Classification
Given the following description of an online meme related to LGBTQ+ movements, analyze: What is the meme’s target subject? Select the most appropriate category from these options: 1. Undirected; 2. Specific Individual; 3. LGBTQ+ Community; 4. Organization. <Target Classification Guidelines> Description of the meme content:[M2T][CoT]
C.4 CoT+: CoT Prompting with Guidelines
The proposed guided CoT reasoning prompts are listed below, where the interpretable human-crafted task-specific guidelines are denoted by <GL>. Detailed guidelines are released in our code.
C.4.1 FHM
Prompt
Given the following image-caption content, which may or may not be an online meme, classify the content as either hateful or non-hateful based on common values, knowledge and social norms.<GL>Image-caption content you need to classify:[M2T].The caption overlaid on the image reads:"[OCR]".[CoT]
Guidelines
@@ C.4.1 FHM
  • Hateful examples generated from Section C.3.
C.4.2 Political Memes
Prompt
Given the following description of an online meme related to {HarMeme: COVID-19 pandemic; Harm-P: U.S. politics; MultiOFF: 2016 U.S. Presidential Election}, classify the meme content as either {HarMeme and Harm-P: harmful or harmless; MultiOFF: offensive or non-offensive} based on widely accepted values, established knowledge and social norms.<GL>Meme content you need to classify:[M2T].[CoT]
Guidelines: HarMeme
@@ C.4.2 Political Memes
C.4.3 MAMI
Prompt
Given the following image-caption content, which may or may not be an online meme, classify the content as either misogynistic or non-misogynistic based on common values, knowledge, social norms and the provided guidelines.<GL>Image-caption content you need to classify:[M2T]. The caption overlaid on the image reads:"[OCR]".[CoT]
Guidelines
@@ C.4.3 MAMI
  • <Caution e.g., Harmful Stereotypes: NOT ALL stereotypes are deemed "harmful". Within the scope of this task, beyond the following provided examples of "harmful" stereotypes against women, other contents that might be interpreted as portraying women in a slightly negative light should not be automatically regarded as "harmful" stereotypes, and therefore should be considered as harmless>, therefore, if the given image-caption content aligns with or intentionally implies any of the following commonly found misogynistic contents, the content should be classified as misogynistic. Such misogynistic contents include: <Examples>. (See our released code for more details.)
@@
  • Unless the content stereotypes women as prone to cheating, avoid overinterpreting content that features dynamics or interactions in "husband-wife" or "boyfriend-girlfriend" relationships to assume negative stereotypes against women.
  • Women vs. men (or boys vs. girls) comparisons are not considered as "harmful" stereotypes against women and should be considered non-misogynistic when such comparison remarks focus on non-sexual daily topics or aspects (e.g., hobbies, interests, attitudes, lifestyles, etc.).
@@
C.4.4 PrideMM
Prompt
Given the following description of an online meme related to LGBTQ+ pride movements, A. If targeting LGBTQ+ community or supporters: classify the content as either harmful or harmless to LGBTQ+ community and supporters, B. If targeting LGBTQ+ individuals: classify the content as either hurtful or non-hurtful to the specific LGBTQ+ individual involved, C. If targeting specific individual: classify the content as either harmful or harmless to the specific individual involved, D. If targeting organizations: classify the content as either harmful or harmless to the public image of the organization(s) involved, according to widely accepted social norms, values, cultural understanding, and the provided guidelines.<GL>Meme content you need to classify:[M2T].[CoT]
Guidelines
@@ C.4.4 PrideMM
  • Try to analyze the meme’s accurate stance toward LGBTQ+ community, supporters and movements. Is the meme’s stance "neutral", "support" or "oppose"?
@@
Appendix D Case Study

### v2

@@ 2. Female Is/Are there any human subject(s) in the image female?[Ignore][Fm]
  1. Politicians (i) Is any politician or celebrity portrayed in the image? If yes, who?[Fm] (ii) Is any head of state portrayed in the image?[Fm] (iii) Is Donald Trump depicted in the image?[Fm] (iv) Is Joe Biden depicted in the image?[Fm] (v) Is Barack Obama depicted in the image?[Fm] (vi) Is Hillary Clinton depicted in the image?[Fm] (vii) Is Bernie Sanders depicted in the image?[Fm] (viii) Is Gary Johnson depicted in the image?[Fm] (ix) Does this image feature Joe Biden and Barack Obama?[Fm]
  2. Political Issues (i) Is any political party explicitly involved in this image?[Fm] (ii) Is LGBTQ+ community or LGBTQ+ individual portrayed in this image?[Fm] (iii) Is any individual of Middle Eastern descent portrayed in the image?[Fm] (iv) Is any protected racial or minority group (such as African Americans or other colored people) portrayed in this image?[Fm]
@@
  1. Describe (i) FHM: What is shown in the image? Describe within two sentences, ignoring any overlaid text or caption. (ii) HarMeme and Harm-P: What is shown in the meme? (iii) MultiOff: What is shown in this image? (iv) MAMI: The overlaid text on the image reads:[OCR]. Question: What is shown in the image? Describe within three sentences. DO NOT assume the nature of the image’s tone or intent as humorous, comical, playful or lighthearted in your description. (v) PrideMM: This is an online meme related to LGBTQ+ pride movement. What is this meme about? Note: DO NOT ASSUME the nature of the meme’s tone and intent as humorous or lighthearted. Describe in a neutral tone.
C.1.2. Information Integration
After acquiring visual information leveraging LMMs, we prompt unimodal text-only LLMs to integrate the extracted visual cues (along with the overlaid caption [OCR] for some datasets) into a coherent, unified description of the content of each meme. We experiment with four open-source small-scale LLMs with no more than 14 billion parameters i.e., Qwen2.5-14B (Report, 2025), Mistral-12B (Mistral, 2024), Qwen2.5-7B (Report, 2025) and Llama3.1-8B (AI@Meta, 2024). All LLMs are the instruction-tuned versions. The visual information gathered is denoted by [VIG]. Examples of integration prompt templates:
  1. FHM Given the following information provided about an image, and disregarding any information about overlaid text or captions, synthesize and rephrase these information into a coherent and unified description of the image’s content. Information:[VIG]
  2. PrideMM Given the following information provided about an online meme, synthesize and rephrase these information into a unified, coherent, and neutral description of the meme content. DO NOT mention the meme’s tone and intent as humorous or light-hearted in the description. Information:[VIG]. The overlaid caption or text recognized in the meme reads:[OCR]
Section C.1.2 converts memes to text. The following sections (including the proposed guided CoT reasoning) are based on the generated meme description, denoted by [M2T]. Chain-of-Thought prompting instruction (e.g., “Now, let’s think step by step:”, “Now, let’s analyze by applying all the guidelines one by one:”) is denoted by [CoT].
C.1.3. Fine-grained Target Identification
@@
  1. Protected Group Detection Given the following image-caption content which may or may not be an online meme, analyze: what vulnerable protected group(s) might be involved in the content? Here are some guidelines for your reference: A. If the content does not appear to involve any specific protected group, just output“No specific protected group involved.” B. If the image content does involve specific protected group(s), choose your answer(s) from the specified list (you may choose multiple options if there are more than one protected groups involved): 1. Women (Female); 2. LGBTQ Community; 3. People with Disabilities; 4. Muslims and Islamic culture; 5. Individuals of Middle Eastern descent; 6. Jewish individuals; 7. Individuals of African descent; 8. African Americans; 9. Individuals of East Asian descent; 10. Individuals of South Asian descent; 11. Native Americans; 12. Other protected groups not listed. Here is the image-caption content you need to analyze:[M2T]. The caption overlaid on the image reads[OCR].[CoT]
We denote the target protected groups identified as $TG_{1}$, $TG_{2}$$TG_{n}$. The following template prompts LLMs with few-shot human-defined hateful examples [FS] to sequentially generate more hateful forms targeting each protected group:
  1. Hateful Forms Generation 1. Provide examples of commonly found harmful stereotypes and forms of offensive, hateful content against $TG_{1}$ in online memes. Provide only phrases or terms without detailed explanations e.g.,[FS]. 2. Provide examples of commonly found …against $TG_{2}$
@@
  1. Entity Identification Given the following description of an online meme related to LGBTQ+ movements, { (i) Country and Region: does the meme explicitly reference any country or region where LGBTQ+ identities or advocacy are discouraged? (ii) Politics: does the meme explicitly involve or mention politicians, political figures, political parties, ideologies, or groups? (iii) Company: does the meme explicitly touch on topics about corporate involvement in LGBTQ+ movements? (iv) Specific Individual: analyze: Does the meme content involve any specific individual? ¡Definition of Specific Individual¿ (v) Organization: analyze: Does the meme content address or discuss organizational involvement related to LGBTQ+ issues? ¡Definition of Organization¿ (vi) LGBTQ+ Subgroup: analyze: what specific subgroup(s) within the LGBTQ+ community is/are particularly mentioned, referenced or implicitly implied in the meme content? ¡Subgroup Identification Guidelines¿ }**Description of the meme content**: [M2T]{[Fm][CoT]}
  2. Target Classification Given the following description of an online meme related to LGBTQ+ movements, analyze: What is the meme’s target subject? Select the most appropriate category from these options: 1. Undirected; 2. Specific Individual; 3. LGBTQ+ Community; 4. Organization. ¡Target Classification Guidelines¿ Description of the meme content:[M2T][CoT]
C.2. CoT+: CoT Prompting with Guidelines
The proposed guided CoT reasoning prompts are listed below, where the interpretable human-crafted task-specific guidelines are denoted by ¡GL¿. Detailed guidelines are released in our codebase.
C.2.1. FHM
  1. Prompt Given the following image-caption content, which may or may not be an online meme, classify the content as either hateful or non-hateful based on common values, knowledge and social norms.¡GL¿Image-caption content you need to classify:[M2T].The caption overlaid on the image reads:[OCR]”.[CoT]
  2. Guidelines
@@
    • Hateful examples generated from Item .
C.2.2. Political Memes
  1. Prompt Given the following description of an online meme related to {HarMeme: COVID-19 pandemic; Harm-P: U.S. politics; MultiOFF: 2016 U.S. Presidential Election}, classify the meme content as either {HarMeme and Harm-P: harmful or harmless; MultiOFF: offensive or non-offensive} based on widely accepted values, established knowledge and social norms.¡GL¿Meme content you need to classify:[M2T].[CoT]
  2. Guidelines for HarMeme
@@ 4. Guidelines for MultiOFF
C.2.3. MAMI
  1. Prompt Given the following image-caption content, which may or may not be an online meme, classify the content as either misogynistic or non-misogynistic based on common values, knowledge, social norms and the provided guidelines.¡GL¿Image-caption content you need to classify:[M2T]. The caption overlaid on the image reads:[OCR]”.[CoT]
  2. Guidelines
@@
    • ¡Caution e.g., Harmful Stereotypes: NOT ALL stereotypes are deemed ”harmful”. Within the scope of this task, beyond the following provided examples of ”harmful” stereotypes against women, other contents that might be interpreted as portraying women in a slightly negative light should not be automatically regarded as ”harmful” stereotypes, and therefore should be considered as harmless¿, therefore, if the given image-caption content aligns with or intentionally implies any of the following commonly found misogynistic contents, the content should be classified as misogynistic. Such misogynistic contents include: ¡Examples¿. (See our released code for more details.)
@@
    • Unless the content stereotypes women as prone to cheating, avoid overinterpreting content that features dynamics or interactions in ”husband-wife” or ”boyfriend-girlfriend” relationships to assume negative stereotypes against women.
    • Women vs. men (or boys vs. girls) comparisons are not considered as ”harmful” stereotypes against women and should be considered non-misogynistic when such comparison remarks focus on non-sexual daily topics or aspects (e.g., hobbies, interests, attitudes, lifestyles, etc.).
@@
C.2.4. PrideMM
  1. Prompt Given the following description of an online meme related to LGBTQ+ pride movements, A. If targeting LGBTQ+ community or supporters: classify the content as either harmful or harmless to LGBTQ+ community and supporters, B. If targeting LGBTQ+ individuals: classify the content as either hurtful or non-hurtful to the specific LGBTQ+ individual involved, C. If targeting specific individual: classify the content as either harmful or harmless to the specific individual involved, D. If targeting organizations: classify the content as either harmful or harmless to the public image of the organization(s) involved, according to widely accepted social norms, values, cultural understanding, and the provided guidelines.¡GL¿Meme content you need to classify:[M2T].[CoT]
  2. Guidelines
@@
    • Try to analyze the meme’s accurate stance toward LGBTQ+ community, supporters and movements. Is the meme’s stance “neutral”, “support” or “oppose”?
@@
C.3. Guideline Perturbations
We apply three different perturbation strategies to our finalized human-crafted guidelines:
  • Rephrasing. Each rule is reworded with GPT-4o to modify surface phrasing while largely maintaining the overall semantics.
  • Shuffling. Both the order of rules and the pattern examples within rules are randomly shuffled.
  • Modifying. The guidelines are altered by revising, adding, or removing rules, or by replacing with a GPT-4o-generated summary. However, this strategy may compromise the coverage and integrity of the guidelines for some datasets.
Table 4 shows that our method remains generally robust to these perturbations, though performance may occasionally fluctuate, particularly on challenging political meme datasets of MultiOFF and Harm-P, where models operate under high uncertainty, making them more sensitive to randomness and semantic nuances introduced by ambiguity from perturbations.
Figure 5. Our proposed High-fidelity Meme2Text pipeline.
Figure 6. Error type distribution (%).
Table 5. Dataset Statistics.
Datasets Test Train
#Harm #Benign
FHM (Kiela et al., 2020) 490 510 8500
HarMeme (Pramanick et al., 2021a) 124 230 3013
Harm-P (Pramanick et al., 2021b) 171 184 2939
MultiOFF (Suryawanshi et al., 2020) 58 91 445
MAMI (Fersini et al., 2022) 500 500 9000
PrideMM (Shah et al., 2024) 247 260 4328
GoatBench (Hateful) (Lin et al., 2024b) 750 1250
GoatBench (Harmful) 420 589
GoatBench (Offensive) 303 440
GoatBench (Misogynistic) 500 500
Table 6. Prediction ensemble strategies. C.L.: Confidence Level for our human-crafted guidelines.
C.L. Dataset M-CoT U-CoT U-CoT+ Decision
1 0 Harmful
ALL 0 1
1 1
High Harm-C, PrideMM 0 0 Harmless
Medium FHM, 0 0 0 0 1 0 Harmless
MultiOFF 0 0 0 0 0 1
0 0
Low Harm-P, 0 0 0 0 Harmless
MAMI 0 0
  • Note: 0/1: Predictions based on visual details extracted by Qwen2VL-7B.
  • 0/1: Predictions based on visual details extracted by LLaVA1.6-7B.
C.4. Guidelines vs. Few shots
We carefully select a small set of representative memes (10–20) from the holdout training split and use their GPT-4o-generated descriptions to construct a candidate pool of demonstration examples for each dataset. We evaluate four few-shot CoT settings with increasing numbers of shots of 4, 6, 8, and 10, where class-balanced examples are randomly sampled from the pool. Results show that positive performance gains $\bar{\Delta}_{\textit{U-CoT+FS}}$ can be found on some datasets, suggesting that our proposed unimodal harmful meme detection framework, built upon the Meme2Text pipeline, could be generalized to few-shot inference settings. Despite specific cases where the performance gap between guideline-based prompting and few-shot prompting is minimal (e.g., on MAMI and PrideMM), the overall in-context learning ability of lightweight LLMs remains limited in detecting harmful memes conveyed through text, making our human-crafted guidelines a more effective solution.
Appendix D DATASETS
Dataset details are summarized as follows:
  • FHM (Kiela et al., 2020) The Facebook Hateful Memes (FHM) dataset was released in 2021 as a challenge set for multimodal classification, focusing on detecting hate speech in multimodal memes. The dataset was synthetically constructed to examine model performance in distinguishing real hateful memes from harmless ones that include their benign confounders, which are constructed with flipped labels through minimal changes to one of the modalities (either images or overlaid text). The hatefulness of a meme can be determined by either both or only one of the modalities (Wu et al., 2020, 2022, 2023a, 2023b, 2024c, 2024d, 2024b, 2024a). Hateful speech under the context of this dataset is defined as: A direct or indirect attack on people based on characteristics, including ethnicity, race, nationality, immigration status, religion, caste, sex, gender identity, sexual orientation, and disability or disease. We define attack as violent or dehumanizing (comparing people to non-human things, e.g. animals) speech, statements of inferiority, and calls for exclusion or segregation. Mocking hate crime is also considered hate speech.
  • HarMeme (Pramanick et al., 2021a) contains 3,544 online memes related to COVID-19 pandemic. They define “harmful” in a broader way different from “hateful” or “offensive” as the potential to cause “mental abuse, defamation, psycho-physiological injury, proprietary damage, emotional disturbance, and compensated public image”. The targets of these harmful memes can be any individual, organization, community, or the general society. Given the inherently ambiguous nature of such definition, which could be different from general criteria internalized in most LLMs, we find that annotated labels in this dataset are highly biased to favor specific types of memes as harmful, particularly those that may harm the public image of Donald Trump.
  • HarmP (Pramanick et al., 2021b) is a follow-up work of HarMeme but has a shift in attention to a broader collection of U.S. political memes, which are labelled into three categories i.e., “very harmful”, “somewhat harmful” and “not harmful”, following the same annotation procedure in HarMeme. Following prior works (Mei et al., 2025), we incorporate “very” and “somewhat” harmful memes into a unified “harmful” class in the experiments. Unlike the conventional definition of “harmful memes”, which typically focuses on harm directed toward society at large or specific vulnerable groups, Harm-P considers a political meme harmful whenever it has the potential to cause harm to any target, including “an individual, an organization, a community, or society as a whole”. Under this definition, harm encompasses a broad range of impacts, such as “mental abuse, defamation, psycho-physiological injury, socio-economic damage, property damage, emotional distress, and damage to public image”. Yet, the ambiguity and variation in such definition of “harmfulness” may have resulted in unclear criteria and inconsistent annotation in the data. In addition, unlike synthetic memes that usually follow a standard creation format, online political memes can be highly free-form or noisy, e.g., digitally altered, distorted, masked, blurred, filled with hardly identifiable text etc. Such low-quality images pose great challenges on existing lightweight LMMs.
  • MultiOFF (Suryawanshi et al., 2020) is a small collection of political memes similar to those in Harm-P but are mainly related to 2016 U.S. presidential election in terms of context. This datasets highlights a specific taxonomy of multiple types of offensive memes, including “racial abuse”, “attacking minorities”, “personal attacks” and “homophobic abuse”.
  • MAMI (Fersini et al., 2022) specifically focuses on misogynistic memes that cause harm to women. Misogynistic memes are generally more explicit in language and imagery compared to other types of harmful memes.
  • PrideMM (Shah et al., 2024) contains 5,063 online memes in the context of LGBTQ+ pride movements. Each meme is annotated with labels indicating its stance toward the goal of LGBTQ+ movements, whether it is considered “hateful” or “humorous”. Apart from the binary labels for hatefulness, the dataset defines a taxonomy of hate targets: “Undirected”, “Individual”, “Community” and “Organization”. Whether a meme can be regarded as “hateful” varies with respect to the perspective of different targets. Therefore, this dataset also includes a target classification task.
  • GoatBench (Lin et al., 2024b) is a comprehensive meme benchmark consisting of over 6K diverse memes that capture themes such as implicit hate speech, sexism, and cyberbullying, curated from the aforementioned datasets. In this work, we evaluate only the harmfulness-related sub-tasks and exclude the sarcasm task.
Appendix E Baselines
We summarize the methodologies and backbone models of the baselines as follows:
  • LMM-RGCL (Mei et al., 2025) learns hate-aware vision and language representations through a contrastive learning objective applied to a pre-trained CLIP encoder, achieving state-of-the-art performance on the MultiOFF dataset.
  • UMR (Yang et al., 2024) integrates uncertainty-guided modal rebalance with Gaussian-based stochastic embeddings and improved cosine loss on frozen CLIP/ALBEF/BLIP/BLIP-2 backbones to handle modality imbalance in multimodal hate detection.
  • Pro-Cap (Cao et al., 2023a) leverages frozen BLIP-2 for zero-shot VQA to generate probing-based captions, which are then combined with meme texts for training BERT and PromptHate models on hateful meme detection.
  • CapAlign (Ji et al., 2024) uses LLaMA-2-7B as the base model and fine-tunes it with LoRA on instruction-following and synthetic alignment conversations, enabling efficient and aligned caption generation for multimodal tasks.
  • MemeCLIP (Shah et al., 2024) builds on frozen CLIP ViT-L/14 encoders and introduces lightweight linear projections, feature adapters, and a cosine classifier with semantic-aware initialization for multi-aspect meme classification.
  • M3H-p-CoT (Kumari et al., 2024) proposes a multi-hop reasoning framework combining EOR features and meme text to prompt a CLIP-based classifier through hierarchical CoT modules trained with cross-attention and contrastive learning.
  • ExplainHM (Lin et al., 2024a) fine-tunes LLaVA (Liu et al., 2023b) to generate multimodal rationales under debate-style prompts and then trains a T5-based model with multimodal fusion to assess meme harmfulness.
  • IntMeme (Hee and Lee, 2025) uses frozen InstructBLIP and mPLUG-Owl to generate meme interpretations in a zero-shot setting, then trains RoBERTa and FLAVA encoders on meme content and interpretations for hateful meme classification.
  • LoReHM (Huang et al., 2024) adopts an agent-based LMM framework (experimenting on LLaVA-34B and GPT-4o) that leverages their few-shot in-context learning and self-improvement capabilities in multi-round chats for low-resource hateful meme detection.
  • Mod-HATE (Cao et al., 2024) trains a suite of LoRA modules and utilizes few-shot demonstration examples to train a module composer, which assigns weights to the LoRA modules for effective low-resource hateful meme detection.
  • GPT-4o (Hurst et al., 2024) is a cutting-edge multimodal model.
Appendix F LIMITATIONS AND FUTURE WORK
We consider the following limitations that we leave for future work: (i) We align our work with ongoing research developments and plan to extend our investigation to culturally diverse meme datasets (e.g., Mandarin, Hindi) as they become more widely available. (ii) Each application of our guideline-crafting procedure across the seven datasets serves as a test of U-CoT+’s cross-context transferability. We leave evaluation on more domain-specific datasets with diverse, underexplored types of harmful memes to future discussions. (iii) Our human-crafted guidelines offer reproducible, easy-to-follow templates that can shed light on future efforts toward automatic guideline generation using LLMs. (iv) Existing datasets lack detailed visual ground truth. Future work will manually assess the robustness of our Meme2Text pipeline to hallucination.
Appendix G CASE STUDY